OpenStudy (anonymous):

Suppose that 11% of all steel shafts produced by a certain process are nonconforming but can be reworked (rather than having to be scrapped). Consider a random sample of 200 shafts, and let X denote the number among these that are nonconforming and can be reworked. probability x is at most 30? less than 30? between 15 and 25?

OpenStudy (anonymous):

idnt understand yuur question

OpenStudy (dumbcow):

p = 0.11 1-p = .89 n = 200 mean = n*p = 22 std dev = n*sqrt(p(1-p)/n) = 4.425 find z values for standard normal distribution Z = (30-22)/4.425 = 1.808 look up table to find P(Z<1.808) which is same P(x <30) for 2nd part, you need 2 Z values Z1 = (25-22)/4.425 = 0.678 Z2 = (15-22)/4.425 = -1.58 P(-1.58 <Z < 0.678) = P(Z<.678) - P(Z<-1.58)